OpenStudy (anonymous):

This is my last one I promise. 2 over 3 + 5 over y-4 = y+6 over 3y-12. Can anyone help solve this please.

OpenStudy (anonymous):

just stop writing "Term1" over "Term2" it's just wrong, cause u use "over" for binomical terms: just write it ("Term1")/("Term"). Simply use brackets.

OpenStudy (anonymous):

Currently u don't get if it's \[\frac{\frac{2}{3+5}}{y-4}=\frac{y+6}{3y-12}\] or \[\frac{2}{\frac{3+5}{y-4}}=\frac{y+6}{3y-12}\]

OpenStudy (anonymous):

Well im sorry I dont know how to write the equations to please you.

OpenStudy (anonymous):

So which one is it? First or second?